OpenStudy (anonymous):

f'(x) = (x)^(1/2)sin x f(0) = 0 Minimum value of f(x) in [0, 10]?

OpenStudy (anonymous):

Look at he graph to help you decide about the Minimum fo f on [0,10] http://www.wolframalpha.com/input/?i=plot++sqrt%28x%29+sin%28x%29+frm+x%3D0+to+x%3D10

OpenStudy (anonymous):

What do you think the answer is? This question is asked in the CLEP Calculus Study Guide, and I don't think the graph even comes close to the supposed answer...

OpenStudy (anonymous):

It does not seem feasible by setting the derivative equal to zero. I found it numerically. The minimum occurs at x=4.81584 and the minimum is -2.18277

OpenStudy (anonymous):

Is your function \[ \sqrt x \, \sin(x) \]

OpenStudy (anonymous):

Yes. The answer is 6.28 btw.

OpenStudy (anonymous):

Something to do with 2pi probably?

OpenStudy (anonymous):

It is not 6.2 for the function \(\Large \sqrt x \, \sin(x) \) on [0,10]

OpenStudy (anonymous):

The graph tells the whole story.

OpenStudy (anonymous):

Hmm.. fair enough.

OpenStudy (anonymous):

Are you sure you copied the function correctly?

OpenStudy (anonymous):

Yes.. it is \[\sqrt{x}\sin x\]

OpenStudy (anonymous):

How can you evaluate this without using a graph?

OpenStudy (anonymous):

You cannot evaluate is explicitly but numerically http://www.wolframalpha.com/input/?i=find+minimum++sqrt%28x%29+sin%28x%29+frm+x%3D0+to+x%3D10

OpenStudy (anonymous):

Sorry, I thought \(f(x)= \sqrt x \sin(x) \) , I did not see that f'(x) was equal that. Hold on a minute

OpenStudy (anonymous):

Yup, that's the case. No worries.

OpenStudy (anonymous):

If you look at the garph of f'(x), you can see that before x=6,28, that f'(x) is negative and after that f'(x) is positive, which means that f has a minimum at x=6.28

OpenStudy (anonymous):

Here is a complete solution. To study the sign of f'(a) on [0, 10] it is enough to study the sign of sin(x) since \(\sqrt x\) is always positive on [0,10] x 0 Pi 2 pi 10 sin(x) 0 + 0 - 0 + f'(x) 0 + 0 - 0 + f(x) inc Max dec Min inc You can see that the minimum occurs at x = 2pi

OpenStudy (anonymous):

Do you understand it now?

OpenStudy (anonymous):

Oh fantastic. Thanks a bunch.

OpenStudy (anonymous):

Okay, so basically.. if a function is decreasing between a certain range, and if one of the limits is the global maxima, the other has to be the minima?